1. State the slope and y-intercept of the graph of 9 – x = 2y



2. 6. Tim has developed a money-saving strategy that uses the formula S = 450 + 100r .
/1 a. Give a possible interpretation of Tim’s strategy.

/1 b. What do the slope and S-intercept of the graph of the relation represent, based on the interpretation from part a.?
/1 c. Using the interpretation from part a., what will have to happen for Tim to save $1250?

/1 d. If Tim plans to stop his savings plan once he has accumulated $1250, what are the domain and range of this relation?

Answer by Boreal(15235)

9-x=2y;
y=(9/2)-(x/2).
Slope is -1/2
y-intercept is 9/2.
Write the equation with y isolated on one side.
----------------------------
s=450+100r; let r=an extra shift at work.
He will save $450 a year regardless plus $100*the number of extra shifts he works per year.
The slope is money gained from extra shifts worked. The S-intercept is the amount he will save regardless.
To save $1250, he has to work at least 8 extra shifts a year.
-------------------------
the domain is 0<=x<=8
The range is 450<=S<=1250.